Problem 2. Solve the following initial value problem y + 2y + y = 4x2 +e-1, y(0) = 0, y(0) = 0

Please answer question fully
Answer #1

y+29+y = 4x + evt 90 20 y(o)=0 Auxillary en (D?+ 2D+1)y=0 1. Let Dam (m2+2m+1) = 0 (m+1)=0 m =-11 Yc (X) = ((1+(x) fx?texht + 191-х, +, ә (+8 - ht-) (9), [в-го) 91 = + о“ (9, 1 * Хуә - 71 +91 - 18 + 9х2 + 2 (x*5+5 - = (9,6 + E) 1 15 x 2 -\ К-